Lane invested $14,000, part at 15% and part at 2%. If the total interest at the end of the year is $1,060, how much did she invest at 15%?
Answer by checkley75(3666)

.15x+.02(14000-x)=1060
.15x+280-.02x=1060
.15x-.02x=1060-280
.13x=780
x=780/.13
x=6000 answer for the amount invested @ 155.
proof
.15*6000+.02*8000=1060
900+160=1060
1060=1060